Use the given information to find the p-value. Use a 0.05 significance level and state the conclusion about the null hypothesis (reject the null hypothesis or fail to reject the null hypothesis).

a. The test statistic in a right-tailed test is z = 1.00.

b. The test statistic in a test with alternative hypothesis p 6= 5/7 is z = 2.05.

c. The test statistic in a test with alternative hypothesis p < 2/3 in z = −0.45.
Answer #1

a) for right tailed test and given test statistic ; p value =P(Z>1) =0.1587

since p value >0.05 , fail to reject the null hypothesis

b)

for two tailed test and given test statistic ; p value =2*P(Z>2.05) =2*0.0202 =0.0404

since p value <0.05 ,  reject the null hypothesis

c)

for left tailed test and  given test statistic ; p value =P(Z<-0.45)=0.3264

since p value >0.05 , fail to reject the null hypothesis